OpenStudy (anonymous):

HELP! if (x1,y1) and (y1,y2) satisfy xy=k, then x1y1 = y1y2. Find x or y as indicated........... --> question (x,x) and (5,25)

jimthompson5910 (jim_thompson5910):

if xy = k, then the first pair (x,x) would multiply to k and so would the second pair (5,25)

jimthompson5910 (jim_thompson5910):

this means that x*x = 5*25

OpenStudy (anonymous):

Wait, ... still confused

jimthompson5910 (jim_thompson5910):

from here, you need to solve x*x = 5*25 for x

OpenStudy (anonymous):

is that multiplication? like 5x25 ?

jimthompson5910 (jim_thompson5910):

* is multiply, yes

jimthompson5910 (jim_thompson5910):

so 5*25 = 5 times 25 = 125

jimthompson5910 (jim_thompson5910):

x*x = x times x = x^2

jimthompson5910 (jim_thompson5910):

so x*x = 5*25 turns into x^2 = 125

OpenStudy (anonymous):

then what?

jimthompson5910 (jim_thompson5910):

well you need to isolate x fully

OpenStudy (anonymous):

so square root it?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

x=5?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

x^2 square rooted is x, and 125 square rooted is 5?

jimthompson5910 (jim_thompson5910):

the square root of 125 is NOT 5

jimthompson5910 (jim_thompson5910):

you're thinking of 25

jimthompson5910 (jim_thompson5910):

use a calculator or you can simplify \(\large \sqrt{125}\)

OpenStudy (anonymous):

My bad, so the answer is x=25?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

I'll show you how to simplify

OpenStudy (anonymous):

Ok. But do I have to solve for y?

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{125} = \sqrt{25*5}\] \[\large \sqrt{125} = \sqrt{25}*\sqrt{5}\] \[\large \sqrt{125} = 5\sqrt{5}\]

jimthompson5910 (jim_thompson5910):

so because the solutions to something like x^2 = 4 are x = 2 or x = -2, this means that the solutions to x^2 = 125 are \[\large 5\sqrt{5}, \ -5\sqrt{5}\]

OpenStudy (anonymous):

and that's the answer?

jimthompson5910 (jim_thompson5910):

yes, the only two possible values for x

OpenStudy (anonymous):

THANK YOU!!!!!!!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

I have another problem.... if you care to help?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

ok, z varies jointly as x and y and inversely as w. Write the appropriate combined-variation equation, and find z for the given values of x, y and w. ---> problem: z=10 when x=5, y=-2. and w=3; x=8, y=6. and w=-12.

jimthompson5910 (jim_thompson5910):

able to get anywhere with this?

OpenStudy (anonymous):

all I know is the formula, is y=kxz/w? I think

jimthompson5910 (jim_thompson5910):

very good

OpenStudy (anonymous):

ok, so how to plug it in is where im struggling.

jimthompson5910 (jim_thompson5910):

the first part tells you that " z=10 when x=5, y=-2. and w=3" plug all this in, and then solve for k

OpenStudy (anonymous):

-2=(k*5*10)/3 ?

jimthompson5910 (jim_thompson5910):

oh sorry, i misread what you said earlier

jimthompson5910 (jim_thompson5910):

the equation should be z = kxy/w

jimthompson5910 (jim_thompson5910):

sorry about that, try it again

OpenStudy (anonymous):

10=(k*5*-2)/3 ?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now solve for k

OpenStudy (anonymous):

ok, so then it would be 10=-10k/3 ?

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

k=-3 ?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

is that the answer then?

jimthompson5910 (jim_thompson5910):

that answers the first part: the part where you need to find the equation

jimthompson5910 (jim_thompson5910):

the equation z = kxy/w turns into z = -3xy/w

jimthompson5910 (jim_thompson5910):

so the equation is z = -3xy/w and you use this to answer the second part

OpenStudy (anonymous):

z=(-3*8*6)/-12

jimthompson5910 (jim_thompson5910):

good, now evaluate

OpenStudy (anonymous):

z=-12

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

close

OpenStudy (anonymous):

what?

jimthompson5910 (jim_thompson5910):

the answer of -12 is incorrect, but it's close

OpenStudy (anonymous):

positive 12 ?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

z = 12

OpenStudy (anonymous):

so that's the final answer right?

jimthompson5910 (jim_thompson5910):

to the second part, yes

OpenStudy (anonymous):

whats next?

jimthompson5910 (jim_thompson5910):

you're done, you found the equation and the value of z given the values of x,y,w

OpenStudy (anonymous):

so when I input my answer, I just have to write z=12?

jimthompson5910 (jim_thompson5910):

the problem asks you to find the equation, which is z = -3xy/w (this is part 1 of the question) then it asks you to find the value of z when x=8, y=6. and w=-12. The answer to this second part is z = 12

OpenStudy (anonymous):

ok, thank you

jimthompson5910 (jim_thompson5910):

you're welcome